# Reliability-Latency-Rate Tradeoff in Low-Latency Communications with Finite-Blocklength Coding

### v3

@@ III Reliability-latency-rate tradeoff in the finite blocklength regime
In this section, we mainly focus on characterizing the reliability-latency-rate tradeoff in the FBL regime for low-latency communication systems. In Section III-A, we start by deriving the high-SNR slope of EC in the FBL regime with fixed error probability, which indicates the service-rate gain of the low-latency communication systems. Based on the results of the high-SNR slope of EC in the FBL regime, we further generalize the conclusions to characterize the reliability gain by representing error probability as a function of SNR in Section III-B. Finally, the gain-conservation equation, which reveals the tradeoff between the service-rate gain, reliability gain, and real-time gain in the FBL and high-SNR regime, is proposed in Section III-C.
@@ III-A High-SNR Slope of EC in the FBL regime
In this subsection, we focus on characterizing the high-SNR slope of EC in the FBL regime. Specifically, we derive the high-SNR slope of EC with a given $\epsilon$ in the AWGN and Nakagami-$m$ fading channel with finite blocklength coding. The derived conclusions can be seen as a generalization of conclusions in [14]. We then provide the corollaries for systems in Rayleigh fading channels with and without frequency or spatial diversity.
@@
We then provide the discussion for the FBL-SISO system in the AWGN and Nakagami-$m$ fading channel, which covers wired communications, satellite communications, and common scenarios in wireless communications. In Theorem 5, we derive the high-SNR slope of the normalized EC for the FBL-SISO system in the AWGN and Nakagami-$m$ fading channel.
Theorem 5. Under the condition that $N=\Psi(\gamma)$ satisfies Theorem 2 with a given $\epsilon$,
@@
See Appendix D.
In Theorem 5, we obtain the value of $s_{\infty}$ in the FBL-SISO system with the AWGN and Nakagami-$m$ fading channels. We can compare this result with the well-known multiplexing gain [42]. Multiplexing gain is defined as the ratio of the maximum coding rate to $\log_{2}\gamma$ as $\gamma\to+\infty$, which is equal to 1 for the SISO system. The definition of multiplexing gain does not include the influence of latency requirements. As we mentioned in Section II-B, $\varrho=\theta N$ indicates the constraint of DVP, which is one of the key performance indicators of QoS. A larger $\theta N$ indicates a stricter QoS constraint according to Eq. (39). Thus, $s_{\infty}$ can be seen as the revised multiplexing gain, which takes the requirements of DVP into account. Since $s_{\infty}$ considers the gain under the QoS constraint, it should not be larger than the multiplexing gain, as validated in Theorem 5.
The results in Theorem 5 show that with a small $\theta N$, $s_{\infty}$ of the fading channel is independent of the statistical constraints, which is consistent with the analysis in [14] for the infinite-blocklength case in the Rayleigh fading channel. This result indicates that if the requirement of DVP is loose, it does not influence the rate of increase of the QoS-constrained throughput. Furthermore, we find that $s_{\infty}$ of the fading channel can be smaller than 1, which is also consistent with the common understanding that the randomness of the fading channel has a negative influence on the QoS-constrained throughput of queueing systems.
Based on Theorem 5, we can obtain $s_{\infty}$ for many common cases. In Corollary 1, we present the $s_{\infty}$ of Rayleigh fading channels with and without frequency or spatial diversity.
Corollary 1. With the same assumption used in Theorem 5,
@@
In Theorem 5 and Corollary 1, we characterize the high-SNR slope of the normalized effective capacity in the FBL regime with a given $\epsilon$. By comparing the results in Theorem 5 and Corollary 1, we find that in the FBL regime, $s_{\infty}$ of the systems with the frequency or spatial diversity is not lower than that of the SISO system in a Rayleigh fading channel. For the SISO system in a Nakagami-$m$ fading channel, its $s_{\infty}$ is not lower than that of the SISO system in a Rayleigh fading channel with $m\geq 1$, while its $s_{\infty}$ is not larger than that of the SISO system in a Rayleigh fading channel with $m\in[0.5,1)$. Furthermore, we also find that, with a larger $\kappa$ or $m$, the length of the interval in which $s_{\infty}$ is irrelevant to $\theta N$ is larger. According to [33], $m$ measures the ratio of line-of-sight signal power to multipath power, which means $m<1$ corresponds to a worse fading than the Rayleigh fading and $m>1$ corresponds to a less severe fading than the Rayleigh fading. Thus, from the above results, we find that the severity of the fading and the spatial or frequency diversity influence $s_{\infty}$. Less serve fading and larger spatial or frequency diversity help us obtain a larger $s_{\infty}$ with a given QoS constraint and ensure $s_{\infty}$ irrelevant to QoS constraints in a larger $\theta N$ set.
Remark 2: We can also define the $s_{0}$ as the low-SNR slope of the normalized EC in the FBL regime. However, we have to determine the denominator in the definition of $s_{0}$. As far as we know, $\gamma$ is not a proper choice since it can result in $\frac{\Lambda(\gamma)}{\gamma}\to+\infty$ when $\gamma$ approaches 0 by adopting a kind of orthogonal code proposed in Chapter 8.5 of [43]. After determining the proper choice of the denominator, we can obtain the upper bound of $s_{0}$ by considering the infinite blocklength with the help of Shannon’s formula. Performing the above analysis is not trivial, which makes it an important future work.
@@ III-B Reliability Gain in the FBL regime
In this subsection, we further consider the case in which $\epsilon$ is a function of $\gamma$. Based on this assumption, we will refine the conclusions in Theorem 5. The revised results are summarized in Theorem 6.
@@
  • for an FBL-SISO system in an AWGN channel, we have
@@
  • for an FBL-SISO system in a Nakagami-$m$ fading channel, we have
    $\lim_{\gamma\to+\infty}\frac{\Lambda(\gamma)}{\log_{2}\gamma}+\log_{2}e\lim_{\gamma\to+\infty}\frac{Q^{-1}(\epsilon(\gamma))}{\sqrt{N}\log_{2}\gamma}=\min\left\{1,\frac{m}{\theta NT\log_{2}e}\right\},$ (49)
@@
See Appendix E.
In Theorem 6, we proposed the result of the high-SNR asymptotic analysis of the EC, which is related to the error probability. Before characterizing the tradeoff between the service-rate gain, reliability gain, and real-time gain, we need to further discuss the reliability gain to build the link between them. First, we define the reliability gain. Let $\varpi=\sqrt{2\log_{2}e}\lim\limits_{\gamma\to+\infty}\frac{\partial\sqrt{-\log_{2}\epsilon(\gamma)}}{\sqrt{N}\partial\log_{2}\gamma}$ denote the reliability gain. $\varpi$ is related to the decay rate of error probability with the increase of $\log_{2}\gamma$ in the high-SNR regime, which is consistent with [30]. With this definition, we can derive the relationship between reliability gain and the second term on the left-hand side of Eqs. (48) and (49). We summarize this relationship in Lemma 1.
Lemma 1. Given a finite reliability gain $\varpi$, it satisfies
@@
See Appendix F.
Through Lemma 1, we can link $s_{\infty}$ and the reliability gain in Eqs. (48) and (49). In Lemma 1, we mention that $c_{4}\in[0,1]$. The value of $c_{4}$ is determined by the form of $\Psi(\gamma)$. When $\lim\limits_{\gamma\to+\infty}\frac{\log_{2}\gamma}{\sqrt{\Psi(\gamma)}}$ exists and is greater than 0, we have $c_{4}=\frac{1}{2}$. This is because, in this case, we find that $\lim\limits_{\gamma\to+\infty}\frac{\Psi(\gamma)}{\log_{2}\gamma}=\lim\limits_{\gamma\to+\infty}\frac{\partial\sqrt{\Psi(\gamma)}}{\partial\log_{2}\gamma}$ is also finite. Thus, we have
@@
$=$ $\lim_{\gamma\to+\infty}\frac{\Psi(\gamma)}{\log_{2}\gamma}\lim_{\gamma\to+\infty}\frac{\log_{2}\gamma}{\sqrt{\Psi(\gamma)}}=1.$
For $\lim\limits_{\gamma\to+\infty}\frac{\log_{2}\gamma}{\sqrt{\Psi(\gamma)}}=0$ or $\lim\limits_{\gamma\to+\infty}\frac{\log_{2}\gamma}{\sqrt{\Psi(\gamma)}}=+\infty$, $\lim\limits_{\gamma\to+\infty}\frac{\partial\sqrt{\Psi(\gamma)}}{\partial\log_{2}\gamma}\frac{\log_{2}\gamma}{\sqrt{\Psi(\gamma)}}$ can be 0, +$\infty$+ ∞, or lies in $(0,+\infty)$.
@@ III-C Gain-Conservation Equation for Reliability, Real-time, and Service-Rate Gains
In this subsection, we will propose the revised gain-conservation equation for low-latency communication systems. In contrast to [30], the revised term in the gain-conservation equation accounts for the influence of random arrivals. Before presenting the revised gain-conservation equation, we first provide the definition of service-rate and real-time gains.
The service-rate gain is defined as $\zeta=\lim\limits_{\gamma\to+\infty}\frac{\partial{\Lambda(\gamma)}}{\partial{\log_{2}\gamma}}$, which is similar to the definition of the service-rate gain in [30]. Note that EC indicates the maximum constant source rate that can be supported by the given service process to satisfy the statistical queueing requirement [13]. Thus, $\zeta$ indicates the rate of increase of throughput with the increase of $\log_{2}\gamma$ in the high-SNR regime. According to L’Hospital’s rule [44], we have
$s_{\infty}=\lim\limits_{\gamma\to+\infty}\frac{\Lambda(\gamma)}{\log_{2}\gamma}=\lim\limits_{\gamma\to+\infty}\frac{\partial{\Lambda(\gamma)}}{\partial{\log_{2}\gamma}}=\zeta.$ (53)
@@
For the real-time gain, we adopt the same definition as proposed in [30], which is given by $\tau=-\frac{1}{D_{\rm max}}\frac{\partial\log_{2}\delta}{\partial\log_{2}\gamma}$. $\tau$ represents the decay rate of DVP with the increase of $\log_{2}\gamma$ in the high-SNR regime.
Based on Theorem 5, Theorem 6, and Lemma 1, we can then obtain the gain-conversation equation for low-latency communication systems with random arrivals in the FBL regime. The result is summarized in Theorem 7.
Theorem 7. In the FBL regime under the same assumption used in Theorem 6
  • for a SISO system in an AWGN channel, we have
@@
  • for a SISO system in a Nakagami-$m$ fading channel with a finite $\varpi$, we have
@@
    By combing Eqs. (57) and (58), we have
    $(1-\theta N\log_{2}e)\zeta+c_{4}\varpi+\tau=\min\left\{1,\frac{m}{\theta NT\log_{2}e}\right\}.$ (59)
@@
Thus, by taking the derivative on both sides of Eq. (60), we have
$-\frac{1}{D_{\rm max}}\lim_{\gamma\to+\infty}\frac{\partial\log_{2}\delta}{\partial\log_{2}\gamma}=\theta N\log_{2}e\lim_{\gamma\to+\infty}\frac{\partial\Lambda(\gamma)}{\partial\log_{2}\gamma}=\left(\theta N\log_{2}e\right)\zeta.$ (61)
@@
In Theorem 7, we obtain the revised gain-conservation equation for low-latency communication systems. There are some differences between Eq. (59) and Eq. (4) in [30]. First, the parameter $c_{4}$ is 1 in [30]. This is mainly because [30] lacks a discussion on the condition under which normal approximation can be used for high-SNR asymptotic analysis. Second, there is $\theta N$ in Eq. (59), which does not exist in [30]. This is because the arrival in [30] is deterministic. The randomness of the arrival causes a deterioration of the performance of queueing systems with respect to DVP and QoS-constrained throughput.
Moreover, we can obtain the real-time gain based on Theorem 5 and Eq. (58) for Nakagami-$m$ fading channels as
@@
Considering the case in which $T=1$, we can treat $\tau$ as a revised version of diversity gain defined in [42]. The diversity gain is defined as the ratio of the outage probability to $\log_{2}\gamma$ with $\gamma\to+\infty$, which is equal to $m$ for the Nakagami-$m$ fading channel. According to L’Hospital’s rule, $\tau$ can be treated as the ratio of DVP to $\log_{2}\gamma$ with $\gamma\to+\infty$. From Eq. (62), if $1\leq\frac{m}{\theta NT\log_{2}e}$, $\theta N\log_{2}e\leq m$. Thus, $\tau\leq m$. The diversity gain proposed in [42] is an upper bound of $\tau$. This is because
@@
Thus, we obtain that $-\frac{1}{D_{\rm max}}\frac{\log_{2}\delta}{\log_{2}\gamma}$ is not larger than the diversity gain proposed in [42].
Corollary 2. In the FBL regime under the assumption used in Theorem 6
  • for a SISO system in a Rayleigh fading channel, we have
    $\tau=\left(\theta N\log_{2}e\right)\zeta.$ (65)
    By combing Eqs. (64) and (65), we have
    $(1-\theta N\log_{2}e)\zeta+c_{4}\varpi+\tau=\min\left\{1,\frac{1}{\theta NT\log_{2}e}\right\}.$ (66)
  • for a system in Rayleigh fading channels with frequency or spatial diversity $\kappa$, we have
    $\tau=\left(\theta N\log_{2}e\right)\zeta.$ (68)
    By combing Eqs. (67) and (68), we have
    $(1-\theta N\log_{2}e)\zeta+c_{4}\varpi+\tau=\min\left\{1,\frac{\kappa}{\theta NT\log_{2}e}\right\}.$ (69)
    In Eqs. (64), (66), (67), and (69), $c_{4}$ is the same as in Eq. (51).

### v4

@@ III Reliability-latency-rate tradeoff in the finite blocklength regime
In this section, we mainly focus on characterizing the reliability-latency-rate tradeoff in the FBL regime for low-latency communication systems. In Section III-A, we begin by deriving the high-SNR slope of EC in the FBL regime with a fixed error probability, which indicates the service-rate gain of low-latency communication systems. Based on the results of the high-SNR slope of EC in the FBL regime, we further generalize the conclusions to characterize the reliability gain by representing error probability as a function of SNR in Section III-B. Finally, in Section III-C, we propose the gain-conservation equation, which reveals the tradeoff among the service-rate gain, reliability gain, and real-time gain in the FBL and high-SNR regimes.
@@ III-A High-SNR Slope of EC in the FBL regime
In this subsection, we focus on characterizing the high-SNR slope of EC in the FBL regime. Specifically, we derive the high-SNR slope of EC with a given $\epsilon$ in both the AWGN and Nakagami-$m$ fading channels with FBC. The derived conclusions can be viewed as a generalization of the conclusions in [14]. We then provide corollaries for systems in Rayleigh fading channels, considering both scenarios with and without frequency or spatial diversity.
@@
We then provide the discussion for the FBL-SISO system in the AWGN and Nakagami-$m$ fading channels, encompassing wired communications, satellite communications, and other typical scenarios in wireless communications. In Theorem 5, we derive the high-SNR slope of the normalized EC for the FBL-SISO system in both the AWGN and Nakagami-$m$ fading channel.
Theorem 5. Under the condition that $N=\Psi(\gamma)$ satisfies Theorem 2 for a given $\epsilon$,
@@
See Appendix D. ∎
In Theorem 5, we derive the value of $s_{\infty}$ for the FBL-SISO system over AWGN and Nakagami-$m$ fading channels. We can compare this result with the well-known multiplexing gain [42]. Multiplexing gain is defined as the ratio of the maximum coding rate to $\log_{2}\gamma$ as $\gamma\to+\infty$, which is equal to 1 for the SISO system. However, the definition of multiplexing gain does not account for latency requirements. As we mentioned in Section II-B, $\varrho=\theta N$ represents the constraint imposed by DVP, which is one of the key performance indicators of QoS. A larger $\theta N$ indicates a stricter QoS constraint, as outlined in Eq. (40). Consequently, $s_{\infty}$ can be interpreted as a revised multiplexing gain, which takes the requirements of DVP into account. Since $s_{\infty}$ considers the gain under the QoS constraint, it can not be larger than the multiplexing gain, as validated in Theorem 5.
The results in Theorem 5 demonstrate that for small values of $\theta N$, $s_{\infty}$ of the fading channel is independent of QoS constraints, which aligns with the analysis in [14] for the infinite-blocklength case in the Rayleigh fading channel. This result indicates that when the requirement of DVP is loose, it does not influence the increasing rate of the QoS-constrained throughput. Moreover, we find that $s_{\infty}$ of the fading channel can be smaller than 1, which is also consistent with the common understanding that the randomness of fading has a negative influence on the QoS-constrained throughput in queueing systems.
Based on Theorem 5, we can obtain $s_{\infty}$ for several common cases. In Corollary 1, we present the $s_{\infty}$ of Rayleigh fading channels with and without frequency or spatial diversity.
Corollary 1. Under the same assumption used in Theorem 5,
@@
In Theorem 5 and Corollary 1, we characterize the high-SNR slope of the normalized EC in the FBL regime with a given $\epsilon$. Comparing the results in Theorem 5 and Corollary 1, we find that in the FBL regime, $s_{\infty}$ of the systems with frequency or spatial diversity is at least as large as that of the SISO system over a Rayleigh fading channel. For the SISO system over a Nakagami-$m$ fading channel, $s_{\infty}$ is no smaller than that of the SISO system over a Rayleigh fading channel with $m\geq 1$. For $m\in[0.5,1)$, $s_{\infty}$ is not larger than that of the SISO system over a Rayleigh fading channel. Furthermore, we observe that as $\kappa$ or $m$ increases, the length of the $\theta N$ interval, where $s_{\infty}$ is irrelevant to $\theta N$, is larger. According to [34], $m$ represents the ratio of line-of-sight signal power to multipath power, with $m<1$ indicating more severe fading than Rayleigh fading and $m>1$ indicating less severe fading than the Rayleigh fading. In summary, the severity of fading and the degree of spatial or frequency diversity both influence $s_{\infty}$. Less serve fading and larger spatial or frequency diversity lead to larger $s_{\infty}$ under a given QoS constraint and ensure that $s_{\infty}$ remains irrelevant to QoS constraints over a broader range of $\theta N$ values.
Remark 2: We can also define $s_{0}$ as the low-SNR slope of the normalized EC in the FBL regime. However, identifying the appropriate denominator in the definition of $s_{0}$ is crucial. As far as we know, $\gamma$ is not a proper choice since it can result in $\frac{\Lambda(\gamma)}{\gamma}\to+\infty$ when $\gamma$ approaches 0 by adopting a kind of orthogonal code proposed in Chapter 8.5 of [43]. Once a proper denominator is determined, we can obtain the upper bound of $s_{0}$ by considering the infinite blocklength regime by leveraging Shannon’s formula. Performing the above analysis is not trivial, which makes it an important future work.
@@ III-B Reliability Gain in the FBL regime
In this subsection, we extend the analysis to the case where $\epsilon$ is a function of $\gamma$. Under this assumption, we will refine the conclusions in Theorem 5. The updated results are summarized in Theorem 6.
@@
  • for an FBL-SISO system over an AWGN channel, we have
@@
  • for an FBL-SISO system over a Nakagami-$m$ fading channel, we have
    $\lim_{\gamma\to+\infty}\frac{\Lambda(\gamma)}{\log_{2}\gamma}+\log_{2}e$ $\lim_{\gamma\to+\infty}\frac{Q^{-1}(\epsilon(\gamma))}{\sqrt{N}\log_{2}\gamma}$ (50)
    $=\min\left\{1,\frac{m}{\theta NT\log_{2}e}\right\}.$
@@
See Appendix E. ∎
In Theorem 6, we proposed the result of the high-SNR asymptotic analysis of the EC, which is related to the error probability. Before characterizing the tradeoff between the service-rate gain, reliability gain, and real-time gain, we need to first explore the concept of reliability gain to establish the link between these metrics. We define the reliability gain as $\varpi=\sqrt{2\log_{2}e}\lim\limits_{\gamma\to+\infty}\frac{\partial\sqrt{-\log_{2}\epsilon(\gamma)}}{\sqrt{N}\partial\log_{2}\gamma}$. $\varpi$ is related to the decay rate of the error probability with respect to $\log_{2}\gamma$ in the high-SNR regime, which is consistent with [30]. Using this definition, we can relate the reliability gain to the second term on the left-hand side of Eqs. (49) and (50), which is summarized in Lemma 1.
Lemma 1. For a finite reliability gain $\varpi$, the following holds:
@@
See Appendix F. ∎
Through Lemma 1, we establish a link between $s_{\infty}$ and the reliability gain in Eqs. (49) and (50). In Lemma 1, we mention that $c_{4}\in[0,1]$. The value of $c_{4}$ depends on the form of $\Psi(\gamma)$. Specifically, when $\lim\limits_{\gamma\to+\infty}\frac{\log_{2}\gamma}{\sqrt{\Psi(\gamma)}}$ exists and is positive, we have $c_{4}=\frac{1}{2}$. This is due to the fact that in this case, $\lim\limits_{\gamma\to+\infty}\frac{\Psi(\gamma)}{\log_{2}\gamma}=\lim\limits_{\gamma\to+\infty}\frac{\partial\sqrt{\Psi(\gamma)}}{\partial\log_{2}\gamma}$ is finite. Thus, we have
@@
$=$ $\lim_{\gamma\to+\infty}\frac{\Psi(\gamma)}{\log_{2}\gamma}\lim_{\gamma\to+\infty}\frac{\log_{2}\gamma}{\sqrt{\Psi(\gamma)}}$
$=$ $1$
For $\lim\limits_{\gamma\to+\infty}\frac{\log_{2}\gamma}{\sqrt{\Psi(\gamma)}}=0$ or $\lim\limits_{\gamma\to+\infty}\frac{\log_{2}\gamma}{\sqrt{\Psi(\gamma)}}=+\infty$, $\lim\limits_{\gamma\to+\infty}\frac{\partial\sqrt{\Psi(\gamma)}}{\partial\log_{2}\gamma}\frac{\log_{2}\gamma}{\sqrt{\Psi(\gamma)}}$ can be 0, +$\infty$+ ∞, or lie in $(0,+\infty)$.
@@ III-C Gain-Conservation Equation for Reliability, Real-time, and Service-Rate Gains
In this subsection, we will propose revised gain-conservation equations for low-latency communication systems. In contrast to [30], the revised terms in the gain-conservation equation account for the influence of random arrivals. Before presenting revised gain-conservation equations, we first provide the definitions of service-rate and real-time gains.
The service-rate gain is defined as $\zeta=\lim\limits_{\gamma\to+\infty}\frac{\partial{\Lambda(\gamma)}}{\partial{\log_{2}\gamma}}$, which is similar to the definition of the service-rate gain presented in [30]. Note that EC represents the maximum constant source rate that can be supported by the given service process to satisfy the statistical queueing requirements [13]. Thus, $\zeta$ indicates the increasing rate of throughput with respect to $\log_{2}\gamma$ in the high-SNR regime. According to L’Hospital’s rule [44], we have
$s_{\infty}$ $=\lim\limits_{\gamma\to+\infty}\frac{\Lambda(\gamma)}{\log_{2}\gamma}$ (54)
$=\lim\limits_{\gamma\to+\infty}\frac{\partial{\Lambda(\gamma)}}{\partial{\log_{2}\gamma}}$
$=\zeta.$
@@
For the real-time gain, we adopt the same definition as proposed in [30], which is given by $\tau=-\frac{1}{D_{\rm max}}\frac{\partial\log_{2}\delta}{\partial\log_{2}\gamma}$. $\tau$ represents the decay rate of DVP as $\log_{2}\gamma$ increases in the high-SNR regime.
Based on Theorem 5, Theorem 6, and Lemma 1, we can then derive revised gain-conversation equations for low-latency communication systems with random arrivals in the FBL regime. The results are summarized in Theorem 7.
Theorem 7. In the FBL regime, under the same assumptions used in Theorem 6,
  • for a SISO system over an AWGN channel, we have
@@
  • for a SISO system over a Nakagami-$m$ fading channel with a finite $\varpi$, we have
@@
    By combining Eqs. (58) and (59), we obtain
    $(1-\theta N\log_{2}e)\zeta+$ $c_{4}\varpi+\tau$ (60)
    $=\min\left\{1,\frac{m}{\theta NT\log_{2}e}\right\}.$
@@
Next, by taking the derivative on both sides of Eq. (61), we have
$-\frac{1}{D_{\rm max}}\lim_{\gamma\to+\infty}\frac{\partial\log_{2}\delta}{\partial\log_{2}\gamma}$ $=\theta N\log_{2}e\lim_{\gamma\to+\infty}\frac{\partial\Lambda(\gamma)}{\partial\log_{2}\gamma}$ (62)
$=\left(\theta N\log_{2}e\right)\zeta.$
@@
In Theorem 7, we present revised gain-conservation equations for low-latency communication systems. There are some differences between Eq. (60) and Eq. (4) in [30]. First, the parameter $c_{4}$ equals 1 in [30]. This is mainly because [30] lacks a discussion on the conditions under which normal approximation can be applied for high-SNR asymptotic analysis. Second, Eq. (60) includes the term $\theta N$, which is absent in [30]. This discrepancy arises because the arrival in [30] is deterministic. The randomness of the arrival causes a deterioration in the performance of queueing systems with respect to DVP and QoS-constrained throughput.
Moreover, we can derive the real-time gain based on Theorem 5 and Eq. (59) for Nakagami-$m$ fading channels as
@@
Considering the case in which $T=1$, we can interpret $\tau$ as a revised version of diversity gain defined in [42]. The diversity gain is defined as the ratio of the outage probability to $\log_{2}\gamma$ as $\gamma\to+\infty$, which equals $m$ for the Nakagami-$m$ fading channel. According to L’Hospital’s rule, $\tau$ can be treated as the ratio of DVP to $\log_{2}\gamma$ as $\gamma\to+\infty$. From Eq. (63), if $1\leq\frac{m}{\theta NT\log_{2}e}$, then $\theta N\log_{2}e\leq m$. Consequently, we have $\tau\leq m$. The diversity gain proposed in [42] is an upper bound of $\tau$, as demonstrated in Eq. (64), which is presented at the bottom of the next page. Thus, we conclude that $-\frac{1}{D_{\rm max}}\frac{\log_{2}\delta}{\log_{2}\gamma}$ is not larger than the diversity gain proposed in [42].
Corollary 2. In the FBL regime, under the assumptions used in Theorem 6,
  • for a SISO system over a Rayleigh fading channel, we have
    $\tau=\left(\theta N\log_{2}e\right)\zeta.$ (66)
    By combining Eqs. (65) and (66), we have
    $(1-\theta N\log_{2}e)\zeta+$ $c_{4}\varpi+\tau$ (67)
    $=\min\left\{1,\frac{1}{\theta NT\log_{2}e}\right\}.$
  • for a system over Rayleigh fading channels with frequency or spatial diversity $\kappa$, we have
    $\tau=\left(\theta N\log_{2}e\right)\zeta.$ (69)
    By combining Eqs. (68) and (69), we have
    $(1-\theta N\log_{2}e)\zeta+$ $c_{4}\varpi+\tau$ (70)
    $=\min\left\{1,\frac{\kappa}{\theta NT\log_{2}e}\right\}.$
    In Eqs. (65), (67), (68), and (70), $c_{4}$ is the same as in Eq. (52).